OpenStudy (anonymous):

Find the derivative of the function f(x) = cos^2 (x) + tan^2 (x).

OpenStudy (anonymous):

\[f \prime(x)=2\cos x(-\sin x)+...\]

OpenStudy (anonymous):

um how did you get that?

OpenStudy (solomonzelman):

ok. \(\large\color{black}{ \cos^2(x) }\) is same as, \(\large\color{black}{ (\cos x)^2 }\). Then he used the CHain rule.

OpenStudy (solomonzelman):

\(\large\color{black}{ \frac{d}{dx}(\cos x)^2 =2(\cos x)^{2-1}\times(- \sin x) }\)

OpenStudy (solomonzelman):

see?

OpenStudy (anonymous):

so you just use the power rule and trig derivative? and the final answer would just be 2cosx x -sinx?

OpenStudy (solomonzelman):

yes, you use the power rule for the cos(x), as if it was just x. THEN, you multiply times the derivative of the inner part. (The inner part is cos(x), and it's derivative is -sin(x). )

OpenStudy (solomonzelman):

So, \(\large\color{black}{ \frac{d}{dx}(\cos x)^2 =2(\cos x)^{2-1}\times(- \sin x)=-2\cos x \sin x }\)

OpenStudy (anonymous):

and you would do the same thing for tan?

OpenStudy (freckles):

yep you would apply this \[\frac{d}{dx}(f(x))^{n}=n(f(x))^{n-1} \cdot f'(x) \]

OpenStudy (solomonzelman):

then you would need the chain rule for \(\large\color{black}{ \tan^2(x) }\). \(\large\color{black}{ \tan^2(x) =(\tan x)^2}\).

OpenStudy (solomonzelman):

(yes, same thing for tan^2x)

OpenStudy (anonymous):

ok so 2cosx(-sinx) x 2tanx(sec^2x)

OpenStudy (solomonzelman):

you had a PLUS in between (inside the function), not a multipkication sign.

OpenStudy (anonymous):

oh oops haha thank you again!

OpenStudy (solomonzelman):

yes, so your answer (as you know, but just want to put it down), is \(\large\color{black}{ f'(x)=-2\cos x \sin x+2\tan x \sec x }\)

OpenStudy (solomonzelman):

you can re-write as, \(\large\color{black}{ f'(x)=2\tan x \sec x-2\sin x \cos x }\)

OpenStudy (solomonzelman):

What does your teacher require to do with the derivative, anything else you need? like re-write in cosines and sines?`

OpenStudy (anonymous):

that is pretty much all she requires, thank you so much!

OpenStudy (solomonzelman):

yw!